OpenStudy (anonymous):

negative 1 over 2 x + 1 = −x + 8

OpenStudy (anonymous):

\[(2x+1)^{-1}=-x+8\]

OpenStudy (anonymous):

wait ion get iht sorry

OpenStudy (radar):

(-1)/(2x) + 1= -x + 8 multiply each term by 2x should get: -1 + 2x = -x^2 + 8x combine terms and simplify \[x ^{2}-6x-1=0\]

OpenStudy (anonymous):